An airplane takes 3 hours to travel a distance of 1680 miles with the wind. The return trip takes 4 hours against the wind. Find the speed of the plane in still air and the speed of the wind.
Answer by mananth(16949)

Plane speed =x mph
wind speed =y mph
against wind 4 hours
with wind 3 hours

Distance against 1680 miles distance with 1680 miles
t=d/r against wind -
1680 / ( x - y )= 4
4 ( x - y ) = 1680
4 x - 4 y = 1680 ....................1

1680 / ( x + y )= 3
3 ( x + y ) = 1680
3 x + 3 y = 1680 ...............2
Multiply (1) by 3
Multiply (2) by 4
we get y
12 x + -12 y = 5040
12 x + 12 = 6720
24 x = 11760
/ 24
x = 490 mph

plug value of x in (1) y
4 x -4 y = 1680
1960 -4 -1960 = 1680
-4 y = 1680
-4 y = -280 mph
y = 70
Plane speed 490 mph
wind speed 70 mph

m.ananth@hotmail.ca